bigboybilly:

LaShawn solved the equation below to the determine the solution. 3 x - 8 = -x + 4 (x - 2) How many solutions does LaShawn’s equation have? zero one two infinitely many

bigboybilly:

can atleas try

Hoodmemes:

Im thinking first u wanna solve them parenthesis on the left side. Like 3 x - 8 = -x + 4 (x - 2) Multiply (2 times 4) then add (-x to 4) this would equal 3x -8 = 3x -8 (i cant give u a direct answer) So ill say its either D=(infinitely many) or A=(zero) Ima hope this helps homie